A laboratory supervisor in hospital is investigating number of reported on-the-job training accidents over a period of one month. Based on past records, she has derived the following probability distribution for x; where x is the number of reported accidents per month. Find the mean


1
Expert's answer
2022-02-25T05:50:53-0500

Needlestick injuries reported "(x)" : 0, 1, 2, 3, 4, 5

Frequency "(f)" : 8, 10, 2, 5, 3, 7


"mean=\dfrac{\sum _ix_if_i}{\sum _if_i}"

"=\dfrac{1}{8+10+2+5+3+7}\big(0(8)+1(10)+2(2)"


"+3(5)+4(3)+5(7)\big)=\dfrac{76}{35}\approx2.17"